OpenStudy (anonymous):

Which is the correct recursive formula for this sequence? -7, -5.6, -4.2, -2.8, -1.4, .... recursive: an = an – 1 + 1.4, where a1 = –7 OR recursive: an = an – 1 + 1.4n, where a1 = –7 ? I tried plugging the numbers into the equation, but couldn't get any option other than 1.4. Please write it in equation form so I can see. Thanks.

OpenStudy (mathmate):

The first 1: a(n)=a(n-1)+1.4 a(0)=-7 a(1)=-7+1.4=-5.6 a(2)=-5.6+1.4=-4.2 ...